# Dissipative Yao-Lee Spin-Orbital Model: Exact Solvability and $\mathcal{P}\mathcal{T}$ Symmetry Breaking

### v1

@@ V $\mathcal{P}\mathcal{T}$-symmetry breaking and Liouvillian exceptional rings
Interestingly, our dissipative model is $\mathcal{P}\mathcal{T}$ symmetric, according to the criteria by Prosen [prosen_PT_criteria]. We will take $\mathcal{P}$ to be
@@
where that we have implicitly used the fact that $\rho$ is Hermitian. In Appendix. A, we detail the criteria for a Liouvillian to be $\mathcal{P}\mathcal{T}$ symmetric [prosen_PT, prosen_PT_criteria] and prove that our model satisfies the conditions, for the above choice of $\mathcal{P}$ and $\mathcal{T}$. Therefore, under the mapping $i\mathcal{L}\to\mathcal{H}$, the traceless part of the non-Hermitian Hamiltonian, defined as $\mathcal{H}^{\prime}=\mathcal{H}-c\mathbb{I}$ with $c=\frac{\text{Tr}(\mathcal{H})}{\text{Tr}(\mathbb{I})}$, is $\mathcal{P}\mathcal{T}$ symmetric in the usual sense for a non-Hermitian Hamiltonian.
@@
We then implement the Fourier transformation by introducing $f_{j\in A}=\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{j}}a_{\bm{k}}$ and $f_{j\in B}=\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{j}}b_{\bm{k}}$ for sublattices $A$ and $B$ respectively, with analogous expressions for the other fermion species $\tilde{f}$. In momentum space, $\mathcal{H}^{\prime}$ can be written as
@@
Diagonalizing Eq. (41), we obtain two doubly degenerate bands,
@@
The eigenvalues $E_{\pm}(\bm{k})$ are purely real when $J^{2}|\Delta(\bm{k})|^{2}>16\gamma^{2}$, corresponding to $\mathcal{P}\mathcal{T}$ preserving modes and purely imaginary Liouvillian eigenvalues $\lambda(\bm{k})$. When $J^{2}|\Delta(\bm{k})|^{2}<16\gamma^{2}$, $E_{\pm}(\bm{k})$ are imaginary and the corresponding Liouvillian modes are decaying in time. The critical points $(k_{x}^{*},k_{y}^{*})$ at which $E=0$ define the boundary between these two qualitatively distinct regimes. At the critical points, the spectrum exhibits a singularity and eigenvalues coalesce in pair, rendering them exceptional points (EPs) of the model [exceptional_spin_liquid, Ding_2022, nonhermitian_topo_liangfu, RevModPhys.93.015005, PhysRevResearch.4.L042025, Özdemir2019]. The location of EPs can be solved as
@@
Note that the $\mathcal{P}\mathcal{T}$ symmetry breaking threshold is different for modes labelled by different momenta. In particular, let $\Delta_{\text{max}}=\max_{\bm{k}}\Delta(\bm{k})=3$ denote the maximum of the structure factor over the Brillouin zone. For $\gamma>\gamma^{*}=\sqrt{J^{2}|\Delta_{\text{max}}(\bm{k})|^{2}/16}=3J/4$, all eigenvalues become entirely imaginary. The momentum-resolved $\mathcal{P}\mathcal{T}$ symmetry breaking is similar to the exceptional rings emerging from interactions and disorders in a Kitaev spin liquid [exceptional_spin_liquid]. However, the non-Hermitian Hamiltonian in our model arises from an exact mapping from the Liouvillian, as opposed to an effective description incorporating effects of the interactions. Therefore, the dissipation strength $\gamma$ can serve as a tuning parameter for the shape and size of the exceptional rings.
@@
To corroborate our analysis in momentum space, we compute the Liouvillian spectrum $\lambda_{n}$ in real space, for different dissipation strengths $\gamma=0.01,0.5,$ and $0.8$. We present the spectrum for a $10\times 10$ lattice in Fig. 4(a). When there is a small but finite dissipation $\gamma<\gamma_{\text{PT}}$, the eigenvalues of $\mathcal{H}^{\prime}$ are all real, and therefore $\lambda_{n}$’s are purely imaginary, corresponding to $\mathcal{P}\mathcal{T}$-preserved modes (left panel of Fig. 4(a)). As $\gamma$ increases above $\gamma_{\text{PT}}$, some modes become $\mathcal{P}\mathcal{T}$-broken and associate with complex $E_{\pm}$, while other modes (within the exceptional rings) remain eigenstates of $\mathcal{P}\mathcal{T}$ and correspond to real $E_{\pm}$. The $\lambda_{n}$’s are therefore a mixture of real and imaginary values (middle panel of Fig. 4(a)). Finally, for $\gamma>\gamma^{*}=3J/4$, all $E_{\pm}(\bm{k})$ become imaginary, leading to an entirely real Liouvillian spectrum (right panel of Fig. 4(a)). Each mode therefore decays with a distinct rate, i.e. are all $\mathcal{P}\mathcal{T}$ broken. Our finding of a three-stage crossover as dissipation strength is turned up has also been analyzed in dissipative fermionic settings [dissipative_stark, quantum_quench_drivendissipative].
@@
Figure 4: (a) Real-space Liouvillian spectrum for $\gamma=0.01$, $\gamma=0.5$ and $\gamma=0.8$. All eigenvalues are shifted by a constant $\text{Tr}(\mathcal{L})/\text{Tr}(\mathbb{I})$ for plotting purposes. The real-space lattice size is $10\times 10$. (b) Estimated $\gamma_{\text{PT}}$ and $\gamma^{*}$ for systems with different linear sizes $L$. The total number of sites is $N=L^{2}$. $\gamma^{*}$ approximately remains constant at $3J/4$ for all system sizes, while $\gamma_{\text{PT}}$ decays as $1/L$, as predicted in Ref. [prosen_PT]. Parameters: $L=4,8,16,32,64;J=1$. (c) Fraction of modes that are $\mathcal{P}\mathcal{T}$ broken. The ratio smoothly crosses over from 0 for $\gamma<\gamma_{\text{PT}}$ to 11 for $\gamma>\gamma^{*}$. Parameters: $L=20$, $J=1$.
@@
The two thresholds, $\gamma_{\text{PT}}$ and $\gamma^{*}$, correspond to dissipation thresholds above which some modes become $\mathcal{P}\mathcal{T}$ broken and all modes are $\mathcal{P}\mathcal{T}$ broken, respectively. $\gamma^{*}$ is only related to the other energy scale in the system, hopping strength $J$, and the lattice geometry. In contrast, $\gamma_{\text{PT}}$ should depend on the system size $N$. We see from Eq. (43) that $\gamma_{\text{PT}}$ is proportional to the minimum of $\Delta(\bm{k})$ over the (discrete) Brillouin zone, $\gamma_{\text{PT}}=J/4\min_{\bm{k}}\Delta(\bm{k})$. In the thermodynamic limit, $\bm{k}$ becomes continuous, and $\min_{\bm{k}}\Delta(\bm{k})\rightarrow\Delta(K)=0$, therefore pushing $\gamma_{\text{PT}}\rightarrow 0$. In finite systems, however, there is usually a small but finite $\gamma_{\text{PT}}$. In Fig. 4(b), we examine the scaling of $\gamma_{\text{PT}}$ and $\gamma^{*}$ with system size. Our numerical results indicate that $\gamma^{*}$ saturates at $3J/4$, while $\gamma_{\text{PT}}$ decreases with the system size, consistent with the inverse level-spacing scaling presented in Ref. [prosen_PT].
@@
To this end, we first need to prepare an initial density matrix $\rho_{0}$ that lies in the flux sector of interest. This can be achieved by applying projectors that enforce the flux configuration. In particular, since $W_{p}$, $\widetilde{W}_{p}$, and $V_{p}$ mutually commute, we can define projectors
$P_{W}$ $=\prod_{p}\frac{1+W_{p}}{2}\cdot\frac{1+\widetilde{W}_{p}}{2};$ (45)
$P_{V}$ $=\prod_{p}\frac{1+V_{p}}{2},$ (46)
which project onto the subspace in which all hexagonal fluxes on both layers and all verticle plaquettes satisfy $W_{p}=\widetilde{W}_{p}=1$, $V_{p}=1$. Then we can define the (properly normalized) initial density matrix as:
$\rho_{0}=\frac{P_{W}P_{V}(\rho)}{\text{Tr}(P_{W}P_{V}(\rho))},$ (47)
where $\rho$ is any initial density matrix. It is easy to check that $W_{p}\rho_{0}=\rho_{0}W_{p}=V_{p}\rho_{0}=\rho_{0}$. In practice, one can choose $\rho$ to be a simple translation-invariant state, such as the maximally mixed state $\rho=\mathbb{I}$ or a pure state $|\rho=\ket{\psi}\!\bra{\psi}$.
Since the initial density matrix is prepared in this sector, and both strong and weak fluxes are conserved quantities, the dynamics remains confined to the sector. Therefore, the Bloch Hamiltonian (Eq. (41)) governs the dynamics of observables and the exceptional ring structure has a tangible consequence on the relaxation dynamics. For example, consider the observable $M=\sum_{i}\sigma_{i}^{y}$, the magnetization in the $y$-direction. Under Liouvillian dynamics, the expectation value of $M$ is:
$\braket{M(t)}=\text{Tr}(M\rho(t))=\text{Tr}(Me^{\mathcal{L}t}\rho_{0}).$ (48)
@@
$\braket{M(t)}$ $=\langle\!\langle M\otimes\mathbb{I}|e^{-i\mathcal{H}t}|\rho_{0}\rangle\!\rangle.$ (49)
In terms of complex fermions, the magnetization can be written as $M=\sum_{i}f_{i}^{\dagger}f_{i}-N,$ i.e. it is the fermion occupation number on the original layer, up to a constant. Note that while the total fermion number is conserved, the number on each layer can change due to the hopping term in Eq. (16).
We can extract the qualitative behavior of this expectation value by computing in momentum space. Importantly, since both $\rho_{0}$ and $M$ are translation-invariant, only the uniform $(\bm{k}=\Gamma=(0,0))$ mode of the Bloch Hamiltonian contributes to $\braket{M(t)}$. All other momenta average to zero when summed over the Brillouin zone. From the spectrum Eq. (43) and the fact that $\Delta(\bm{k}=(0,0))=3$, we see that for $\gamma<3J/4$, the eigenvalues of $h(\bm{k})$ are entirely real, corresponding to oscillatory dynamics. On the other hand, for $\gamma>3J/4$, $E_{\pm}(\bm{k}=(0,0))$ becomes imaginary, leading to decaying Liouvillian modes. Finally, reinstating the constant shift of Liouvillian spectrum (which we denote as $c_{L}=\frac{\text{Tr}(\mathcal{L})}{\text{Tr}(\mathbb{I})}$ and manifests as a rescaling of the observables [quantum_quench_drivendissipative]), we expect the magnetization to scale as:
$\braket{M(t)}\sim e^{-c_{L}t}\times\left\{\begin{array}[]{l}\cos(\sqrt{9J^{2}-16\gamma^{2}}t),\gamma<\frac{3}{4}J;\\ e^{-\sqrt{16\gamma^{2}-9J^{2}}t},\gamma>\frac{3}{4}J,\end{array}\right$. (50)
There is a clear, qualitative difference in the dynamics. Specifically, up to a decaying envelope, $\braket{M(t)}$ remains oscillatory throughout the $\mathcal{P}\mathcal{T}$ mixed phase and transitions to decaying when $\gamma>\gamma^{*}$. This could in principle be observed in experimental platforms where both the spin-orbital Hamiltonian and local spin dissipations could be engineered. To probe other points in momentum space, one could measure spatially modulated observables.

### v2

@@ V $\mathcal{P}\mathcal{T}$-symmetry breaking and Liouvillian exceptional rings
Interestingly, our dissipative model is $\mathcal{P}\mathcal{T}$ symmetric, according to the criteria by Prosen [62]. We will take $\mathcal{P}$ to be
@@
where that we have implicitly used the fact that $\rho$ is Hermitian. In Appendix. A, we detail the criteria for a Liouvillian to be $\mathcal{P}\mathcal{T}$ symmetric [63, 62] and prove that our model satisfies the conditions, for the above choice of $\mathcal{P}$ and $\mathcal{T}$. Therefore, under the mapping $i\mathcal{L}\to\mathcal{H}$, the traceless part of the non-Hermitian Hamiltonian, defined as $\mathcal{H}^{\prime}=\mathcal{H}-c\mathbb{I}$ with $c=\frac{\text{Tr}(\mathcal{H})}{\text{Tr}(\mathbb{I})}$, is $\mathcal{P}\mathcal{T}$ symmetric in the usual sense for a non-Hermitian Hamiltonian.
@@
We then implement the Fourier transformation by introducing $f_{j\in A}=\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{j}}a_{\bm{k}}$ and $f_{j\in B}=\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{j}}b_{\bm{k}}$ for sublattices AA and $B$ respectively, with analogous expressions for the other fermion species $\tilde{f}$. In momentum space, $\mathcal{H}^{\prime}$ can be written as
@@
We note that the $\mathcal{P}\mathcal{T}$ symmetry preserves the Bloch Hamiltonian $h(\bm{k})$. Since $\mathcal{P}$ is implemented through a parity operator, and $\mathcal{T}$ is complex conjugation, $\mathcal{P}\mathcal{T}$ acts on the Bloch Hamiltonian as:
$(\mathcal{P}\mathcal{T})h(\mathbf{k})(\mathcal{P}\mathcal{T})^{-1}=h^{*}(\mathbf{-k})=h(\mathbf{k}),$ (43)
where we have used the fact that $\Delta(-\mathbf{k})=\Delta^{*}(\mathbf{k}).$
Diagonalizing the Bloch Hamiltonian (Eq. (41)), we obtain two doubly degenerate bands,
@@
The eigenvalues $E_{\pm}(\bm{k})$ are purely real when $J^{2}|\Delta(\bm{k})|^{2}>16\gamma^{2}$, corresponding to $\mathcal{P}\mathcal{T}$ preserving modes and purely imaginary Liouvillian eigenvalues $\lambda(\bm{k})$. When $J^{2}|\Delta(\bm{k})|^{2}<16\gamma^{2}$, $E_{\pm}(\bm{k})$ are imaginary and the corresponding Liouvillian modes are decaying in time. The critical points $(k_{x}^{*},k_{y}^{*})$ at which $E=0$ define the boundary between these two qualitatively distinct regimes. At the critical points, the spectrum exhibits a singularity and eigenvalues coalesce in pair, rendering them exceptional points (EPs) of the model [85, 22, 38, 8, 86, Özdemir2019]. The location of EPs can be solved as
@@
Note that the $\mathcal{P}\mathcal{T}$ symmetry breaking threshold is different for modes labelled by different momenta. In particular, let $\Delta_{\text{max}}=\max_{\bm{k}}\Delta(\bm{k})=3$ denote the maximum of the structure factor over the Brillouin zone. For $\gamma>\gamma^{*}=\sqrt{J^{2}|\Delta_{\text{max}}(\bm{k})|^{2}/16}=3J/4$, all eigenvalues become entirely imaginary. The momentum-resolved $\mathcal{P}\mathcal{T}$ symmetry breaking is similar to the exceptional rings emerging from interactions and disorders in a Kitaev spin liquid [85]. However, the non-Hermitian Hamiltonian in our model arises from an exact mapping from the Liouvillian, as opposed to an effective description incorporating effects of the interactions. Therefore, the dissipation strength $\gamma$ can serve as a tuning parameter for the shape and size of the exceptional rings.
@@
To corroborate our analysis in momentum space, we compute the Liouvillian spectrum $\lambda_{n}$ in real space, for different dissipation strengths $\gamma=0.01,0.5,$ and $0.8$. We present the spectrum for a $10\times 10$ lattice in Fig. 4(a). When there is a small but finite dissipation $\gamma<\gamma_{\text{PT}}$, the eigenvalues of $\mathcal{H}^{\prime}$ are all real, and therefore $\lambda_{n}$’s are purely imaginary, corresponding to $\mathcal{P}\mathcal{T}$-preserved modes (left panel of Fig. 4(a)). As $\gamma$ increases above $\gamma_{\text{PT}}$, some modes become $\mathcal{P}\mathcal{T}$-broken and associate with complex $E_{\pm}$, while other modes (within the exceptional rings) remain eigenstates of $\mathcal{P}\mathcal{T}$ and correspond to real $E_{\pm}$. The $\lambda_{n}$’s are therefore a mixture of real and imaginary values (middle panel of Fig. 4(a)). Finally, for $\gamma>\gamma^{*}=3J/4$, all $E_{\pm}(\bm{k})$ become imaginary, leading to an entirely real Liouvillian spectrum (right panel of Fig. 4(a)). Each mode therefore decays with a distinct rate, i.e. are all $\mathcal{P}\mathcal{T}$ broken. Our finding of a three-stage crossover as dissipation strength is turned up has also been analyzed in dissipative fermionic settings [89, 78].
@@
Figure 4: (a) Real-space Liouvillian spectrum for $\gamma=0.01$, $\gamma=0.5$ and $\gamma=0.8$. All eigenvalues are shifted by a constant $\text{Tr}(\mathcal{L})/\text{Tr}(\mathbb{I})$ for plotting purposes. The real-space lattice size is $10\times 10$. (b) Estimated $\gamma_{\text{PT}}$ and $\gamma^{*}$ for systems with different linear sizes $L$. The total number of sites is $N=L^{2}$. $\gamma^{*}$ approximately remains constant at $3J/4$ for all system sizes, while $\gamma_{\text{PT}}$ decays as $1/L$, as predicted in Ref. [63]. Parameters: $L=4,8,16,32,64;J=1$. (c) Fraction of modes that are $\mathcal{P}\mathcal{T}$ broken. The ratio smoothly crosses over from 0 for $\gamma<\gamma_{\text{PT}}$ to 11 for $\gamma>\gamma^{*}$. Parameters: $L=20$, $J=1$.
@@
The two thresholds, $\gamma_{\text{PT}}$ and $\gamma^{*}$, correspond to dissipation thresholds above which some modes become $\mathcal{P}\mathcal{T}$ broken and all modes are $\mathcal{P}\mathcal{T}$ broken, respectively. $\gamma^{*}$ is only related to the other energy scale in the system, hopping strength $J$, and the lattice geometry. In contrast, $\gamma_{\text{PT}}$ should depend on the system size NN. We see from Eq. (44) that $\gamma_{\text{PT}}$ is proportional to the minimum of $\Delta(\bm{k})$ over the (discrete) Brillouin zone, $\gamma_{\text{PT}}=J/4\min_{\bm{k}}\Delta(\bm{k})$. In the thermodynamic limit, $\bm{k}$ becomes continuous, and $\min_{\bm{k}}\Delta(\bm{k})\rightarrow\Delta(K)=0$, therefore pushing $\gamma_{\text{PT}}\rightarrow 0$. In finite systems, however, there is usually a small but finite $\gamma_{\text{PT}}$. In Fig. 4(b), we examine the scaling of $\gamma_{\text{PT}}$ and $\gamma^{*}$ with system size. Our numerical results indicate that $\gamma^{*}$ saturates at $3J/4$, while $\gamma_{\text{PT}}$ decreases with the system size, consistent with the inverse level-spacing scaling presented in Ref. [63].
@@
To this end, we first need to prepare an initial density matrix $\rho_{0}$ that partially lies in the flux sector of interest. Note that no physical density matrix can live entirely in the $V_{p}=1$ sector, due to its traceless-ness. Concretely, we consider two product states, defined as:
[…]
and construct the following density matrix:
[…]
where
[…]
From the definition of the weak $\mathbb{Z}_{2}$ symmetry in terms of the spin variables, Eq. (26), one can easily check that $\rho_{+}$ has eigenvalue +1+1 under the $\mathbb{Z}_{2}$ weak symmetry and thus lives entirely in the $W_{p}=\widetilde{W}_{p}=1;V_{p}=-1$ sector, while $\rho_{-}$ falls in the $W_{p}=\widetilde{W}_{p}=V_{p}=1$ sector. $\rho_{+}$ describes the “population”, or “coherence”, between the two states, while $\rho_{-}$ can be interpreted as the interference. Note that $\text{Tr}(\rho_{+})=1$, while $\text{Tr}(\rho_{-})=0$, which is consistent with the observation in Sec. IV.2 that only $V_{p}=-1$ hosts steady states, which are of trace 1.
Next we consider a specific observable $M=\sum_{i}\sigma_{i}^{y}$, the magnetization in the $y$-direction. In terms of complex fermions, the magnetization can be written as $M=\sum_{i}f_{i}^{\dagger}f_{i}-N,$ i.e. it is the fermion occupation number on the original layer, up to a constant. Note that while the total fermion number is conserved, the number on each layer can change due to the hopping term in Eq. (16).
Under Liouvillian dynamics, the expectation value of $M$ is:
$\braket{M(t)}$ $=\text{Tr}(M\rho(t))=\text{Tr}(Me^{\mathcal{L}t}\rho_{0})$ (52)
$=\text{Tr}(Me^{\mathcal{L}t}\rho_{+})+\text{Tr}(Me^{\mathcal{L}t}\rho_{-}).$ (53)
@@
$\braket{M(t)}$ $=\langle\!\langle M\otimes\mathbb{I}|e^{-i\mathcal{H}t}|\rho_{+}\rangle\!\rangle+\langle\!\langle M\otimes\mathbb{I}|e^{-i\mathcal{H}t}|\rho_{-}\rangle\!\rangle.$ (54)
Crucially, since both strong and weak fluxes are conserved quantities, the dynamics remains confined to each sector. This fact allows for computing the two terms separately, as their dynamics are independent. We can extract the qualitative behavior of the expectation value by working in momentum space. Importantly, since both $\rho_{+}$, $\rho_{-}$, and $M$ are translation-invariant, only the uniform $(\bm{k}=\Gamma=(0,0))$ mode of the Bloch Hamiltonian contributes to $\braket{M(t)}$. All other momenta average to zero when summed over the Brillouin zone.
Let us first consider $\langle\!\langle M\otimes\mathbb{I}|e^{-i\mathcal{H}t}|\rho_{+}\rangle\!\rangle$. Since the desired flux configuration is $W_{p}=\widetilde{W}_{p}=1;V_{p}=-1$, we choose the bond variables as:
$u_{ij}=1;\tilde{u}_{ij}=-1;v_{i}=1\,\,\forall\,i,j.$ (55)
In momentum space, the Bloch Hamiltonian is now
$h(\bm{k})^{V_{p}=-1}=\frac{1}{2}\begin{pmatrix}0&Ji\Delta(\bm{k})&4\gamma&0\\ -Ji\Delta(\bm{k})^{*}&0&0&4\gamma\\ -4\gamma&0&0&Ji\Delta(\bm{k})\\ 0&-4\gamma&-Ji\Delta(\bm{k})^{*}&0\end{pmatrix}$. (56)
Note the difference between the Bloch Hamiltonian in this sector, compared to one in Eq. 41. The spectrum in this sector is thus:
$E_{\pm}^{V_{p}=-1}(\mathbf{k})=-2i\gamma+\left(\pm 2i\gamma\pm\frac{J}{2}|\Delta(\bm{k})|\right).$ (57)
Crucially, the real and imaginary parts of the spectrum decouple, and modes are either oscillatory, or decaying with a momentum-independent rate. Therefore, qualitatively, we expect:
$M_{V_{p}=-1}=\langle\!\langle M\otimes\mathbb{I}|e^{-i\mathcal{H}t}|\rho_{+}\rangle\!\rangle\sim C(M_{0}+e^{-2\gamma t}).$ (58)
Next, we compute $\langle\!\langle M\otimes\mathbb{I}|e^{-i\mathcal{H}t}|\rho_{-}\rangle\!\rangle$. From the spectrum Eq. (44) and the fact that $\Delta(\bm{k}=(0,0))=3$, we see that for $\gamma<3J/4$, the eigenvalues of $h(\bm{k})$ are entirely real, corresponding to oscillatory dynamics. On the other hand, for $\gamma>3J/4$, $E_{\pm}(\bm{k}=(0,0))$ becomes imaginary, leading to decaying Liouvillian modes. Finally, reinstating the constant shift of Liouvillian spectrum (which we denote as $c_{L}=\frac{\text{Tr}(\mathcal{L})}{\text{Tr}(\mathbb{I})}$ and manifests as a rescaling of the observables [78]), we expect the magnetization to scale as:
$\braket{M(t)}\sim M_{V_{p}=-1}+e^{-c_{L}t}\times\left\{\begin{array}[]{l}\cos(\sqrt{9J^{2}-16\gamma^{2}}t),\gamma<\frac{3}{4}J;\\ e^{-\sqrt{16\gamma^{2}-9J^{2}}t},\gamma>\frac{3}{4}J,\end{array}\right$. (59)
There is a clear, qualitative difference in the dynamics. Specifically, up to a term from the steady state sector and an envelope, $\braket{M(t)}$ remains oscillatory throughout the $\mathcal{P}\mathcal{T}$ mixed phase and transitions to decaying when $\gamma>\gamma^{*}$. This could in principle be observed in experimental platforms where both the spin-orbital Hamiltonian and local spin dissipations could be engineered. To probe other points in momentum space, one could measure spatially modulated observables.